Problem with ADAM

the problem found yesterday has been verified to be a problem in the adam_wrapper lib. SOLVED. It is still not understood because this criticality in the software papered just yesterday.

test of 4D trigger

  • tested with i4d_opt_intmat(from=0, to=100): OK (see pictures of the oscilloscope)
  • indirect test of disturbance: OK

Final tests before installing in Hub

Note: we used the last flat available just for increase the forces during the test. Not as verification of flattening. That verification will be done optically next days. Moreover the FF matrix is still the old one (not the one recalibrate yesterday) to check the system in a worse condition

  • 90 deg elevation:
    • shell set and flat: OK. see telemetry from 20100917_163919 to from 20100917_163944
    • noise: max noise 4.6nm rms OK. 20100917_164009
    • step response: max ST=0.96ms. Overshoot < 6%. OK. 20100917_164349
    • turbulence tracking: No done, problem with the IDL software (no restore the disturb file)

  • slewed from 90deg to 20deg elevation with shell set

  • 20 deg elevation:
    • shell rested: gap OK. see telemetry from 20100917_165910 to 20100917_165920
    • setpistoncurrent(20.0)
    • shell set and flat: see telemetry from 20100917_170106 to from 20100917_170137
    • noise: max noise 5.2nm rms OK. 20100917_170151
    • step response: max ST=0.96ms. Overshoot < 6%. OK. 20100917_170403
    • turbulence tracking (if there is time): No done, problem with the IDL software

  • slew elev. test 20-90-20deg several times:
    • see telemetry from 171222 to 172558
    • the difference in force and position is concentrated in the center as expected
    • no reset during slewing, no relevant drift of forces. OK

  • slew elev. test 90-20-90deg several times:
    • set and flat at 90deg elevation
    • for slew see telemetry from 172840 to 172558
    • the difference in force and position is concentrated in the center as expected
    • no reset during slewing, no relevant drift of forces. OK

Setting of the retention ring

  • used 400um shim. 400um exit hard after tightening the screws of the retention ring. Checked that the 300um shim goes loose in the gap. The gap is everywhere larger of 300um for sure.
  • Kapton tape on the joints between miniskirt sectors to seal

slewing test for checking retention ring

  • slew elev. test 90-20deg: no interference with retention ring. See telemetry from 200340 to200438
  • shell ripped by user at 20deg el. and checked the gap all around that the 300um shim goes loose in the gap
  • shell set at 20deg el.: see telemetry form 201700 to 201720

Installation of unit in the HUB

  • XXX

-- ArmandoRiccardi - 17 Sep 2010
I Attachment Action Size Date Who Comment
tek00000.pngpng tek00000.png manage 13 K 17 Sep 2010 - 18:40 ArmandoRiccardi Spare Switch BCU: 4D trigger generation
tek00001.pngpng tek00001.png manage 20 K 17 Sep 2010 - 18:40 ArmandoRiccardi Spare Switch BCU: 4D trigger generation
Topic revision: r2 - 17 Sep 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback